You work for an insurance company that processes claims through both a newer, larger high-tech facility and an older, smaller low-tech facility. Each month, the high-tech facility is working at capacity handling 10,000 claims. It incurs $100,000 in fixed costs and $100,000 in variable costs (which is almost entirely labor costs). Each month, the low-tech facility handles 2,000 claims, incurs $16,000 in fixed costs and $24,000 in variable costs (which is almost entirely labor costs). If you anticipate that in the short-run the number of claims will decrease by 1,000 per month, where does it most make sense (from a cost point of view) to reduce claims processing and lay-off workers first?
Answer #1

Solution:

Facility

No.of claims

FC

VC

TC

VC per claim

TC per claim

High-tech

10,000

100,000

100,000

200,000

$100,000/10,000 = 10

$200,000/10,000 = $20

Low-Tech

2,000

16,000

24,000

40,000

$24,000 /2,000 = 12

$24,000 / 2,000 = $20

 

As fixed costs are fixed thus need to be ignored for this problem, and thus consider only the average variable cost for processing the claim at each location. The above calculations reflects that Low-Tech facilty has higher average variable cost, thus from the low tech facilty lay off the workers.